# SuperM2M: Supervised and Mixture-to-Mixture Co-Learning for Speech Enhancement and Robust ASR

### v2

@@ Abstract
Neural speech enhancement , robust ASR
journal: a journal
@@ 1 Introduction
On the other hand, during data collection, besides using far-field microphones to record target speech, a close-talk microphone is often placed near the target speaker to collect its close-talk speech (e.g., in the CHiME [36], AMI [37], AliMeeting [38], and MISP [39] setup). [^1] Close-talk speech is almost always recorded together with far-field speech in speech separation and recognition datasets, as it is much easier for humans to annotate word transcriptions and speaker activities based on close-talk recordings (where the target speech is very strong) than far-field recordings. See Fig. 1 for an illustration. Although the close-talk microphone can also pick up non-target signals, the recorded close-talk mixture typically has a much higher signal-to-noise ratio (SNR) of the target speaker than any far-field mixtures. In our recent mixture to mixture (M2M) algorithm [40], which builds upon UNSSOR, we find that, besides far-field mixtures, the close-talk mixture can also be leveraged as a weak supervision for training DNNs to separate mixed speakers.
Figure 1: Task illustration. Close-talk mixture consists of close-talk speech and non-target signals. Far-field mixture consists of far-field speech and non-target signals. Best in color.
@@
In this context, we propose to extend UNSSOR [35] and M2M [40] for speech enhancement and robust ASR, where we train and evaluate the proposed algorithms not only on simulated data but also on real data. Our major goal is to show whether the resulting algorithms can yield better generalization to real data than purely-supervised models trained on simulated data, a demonstration that is missing in UNSSOR [35] and M2M [40]. Without this demonstration, the evidence supporting whether this un- and weakly-supervised line of research is worth investigating would be lacking, especially considering the dominance and simplicity of purely-supervised approaches based on simulated training data. We summarize the key contributions of this paper as follows:
@@
  • Considering noise sources as a single, combined source, we propose to formulate the training of speech enhancement models on real data as solving a blind deconvolution problem, following the formulations in UNSSOR [35] and M2M [40] designed for speaker separation.
  • We propose SuperM2M, a co-learning strategy which trains the same DNN model by alternating between M2M training on real data and supervised learning on simulated data. This way, M2M can benefit from massive simulated training data, especially when the real training data is scarce. In addition, we find that this strategy can help mitigate the weaknesses of UNSSOR and M2M on source and frequency permutation, and source ambiguity (which we will detail in Section 4.5).
@@
The evaluation results provide an experimental evidence supporting the strong potential of our un- and weakly-supervised line of research for speech enhancement. A sound demo is provided in the link below. [^2] https://zqwang7.github.io/demos/SuperM2M_demo/index.html
@@ 2.1 Frontend Enhancement and Robust ASR
Diffierently, in this paper we aim at building neural speech enhancement models whose immediate estimate itself can have low distortion to target speech and high reduction to non-target signals, especially on real test data. We find that, on the challenging real test data of CHiME-4 [42], the immediate output of SuperM2M bears low distortion to target speech and high reduction to non-target signals, and feeding it directly to strong ASR models for recognition yields strong performance.
@@ 3 Problem Formulation
We starts with describing the hypothesized physical models, then propose to formulate speech enhancement as a blind deconvolution problem, and at last overview the proposed algorithm for speech enhancement.
@@ 3.1 Physical Model
In reverberant conditions with a compact far-field $P$-microphone array and a single target speaker wearing a close-talk microphone (see Fig. 1 for an illustration), the physical model for each recorded mixture can be formulated, in the short-time Fourier transform (STFT) domain, in the following way.
At a designated reference far-field microphone $q\in\{1,\dots,P\}$, the recorded mixture is formulated as
$Y_{q}(t,f)$ $=X_{q}(t,f)+V_{q}(t,f)$
$=S_{q}(t,f)+H_{q}(t,f)+V_{q}(t,f)$
$=S_{q}(t,f)+\mathbf{g}_{q}(f)^{\mathsf{H}}\widetilde{\mathbf{S}}_{q}(t,f)+V_{q}(t,f)+\varepsilon_{q}(t,f),$ (1)
where $t$ indexes $T$ frames, $f$ indexes $F$ frequency bins, and $Y_{q}(t,f)$, $X_{q}(t,f)$ and $V_{q}(t,f)$ in row $1$ are respectively the STFT coefficients of the mixture, reverberant speech of the target speaker, and non-speech signals at time $t$, frequency $f$ and microphone $q$. In row $2$, $X_{q}(t,f)$ is decomposed to direct-path signal $S_{q}(t,f)$ and reverberation $H_{q}(t,f)$. In row $3$, following narrowband approximation [66, 51], reverberation $H_{q}(\cdot,f)$ is approximated as a linear convolution between the direct-path signal $S(\cdot,f)$ and a linear filter $\mathbf{g}_{q}(f)$. That is, $H_{q}(t,f)\approx\mathbf{g}_{q}(f)^{\mathsf{H}}\widetilde{\mathbf{S}}_{q}(t,f)$, where $\widetilde{\mathbf{S}}_{q}(t,f)=[\hat{S}_{q}(t-K+1,f),\dots,\hat{S}_{q}(t-\Delta,f)]\in{\mathbb{C}}^{K-\Delta}$ stacks $K-\Delta$ T-F units with $\Delta$ denoting a positive prediction delay, $\mathbf{g}_{q}(f)\in{\mathbb{C}}^{K-\Delta}$ can be interpreted as the relative transfer function (RTF) relating the direct-path signal to the reverberation of the target speaker, and $(\cdot)^{\mathsf{H}}$ computes Hermitian transpose. In row $3$, $\varepsilon_{q}(\cdot,f)$ denotes the modeling error incurred by narrowband approximation. In the rest of this paper, when dropping indices $t$ and $f$, we refer to the corresponding spectrograms. We emphasize that $V_{q}$ could contain multiple strong, non-stationary directional as well as diffuse noises. In this paper, we model them using a single, combined source.
At any non-reference far-field microphone $p\in\{1,\dots,P\}$, where $p\neq q$, we formulate the physical model as
@@
$=\mathbf{h}_{p}(f)^{\mathsf{H}}\overline{\mathbf{S}}_{q}(t,f)+V_{p}(t,f)+\varepsilon_{p}(t,f)$
$=\mathbf{h}_{p}(f)^{\mathsf{H}}\overline{\mathbf{S}}_{q}(t,f)+\mathbf{r}_{p}(f)^{\mathsf{H}}\acute{\mathbf{V}}_{q}(t,f)+\varepsilon^{\prime}_{p}(t,f).$ (2)
@@
In row $2$, we use narrowband approximation, similarly to (1), to approximate $X_{p}(t,f)\approx\mathbf{h}_{p}(f)^{\mathsf{H}}\overline{\mathbf{S}}_{q}(t,f)$, where $\overline{\mathbf{S}}_{q}(t,f)=[S_{q}(t-\dot{I}+1,f),\dots,S_{q}(t+\dot{J},f)]\in{\mathbb{C}}^{\dot{I}+\dot{J}}$ and $\mathbf{h}_{p}(f)\in{\mathbb{C}}^{\dot{I}+\dot{J}}$, and we use $\varepsilon_{p}$ to denote the modeling error. $\mathbf{h}_{p}(f)$ can be interpreted as the RTF relating the direct-path signal $S_{q}$ (captured by the reference microphone $q$) to the speaker image at another far-field microphone $p$ (i.e., $X_{p}$). In row $3$, we use the same trick to approximate non-speech signals $V_{p}(t,f)\approx\mathbf{r}_{p}(f)^{\mathsf{H}}\acute{\mathbf{V}}_{q}(t,f)$, where $\acute{\mathbf{V}}_{q}(t,f)=[V_{q}(t-\dot{I}+1,f),\dots,V_{q}(t+\dot{J},f)]\in{\mathbb{C}}^{\dot{I}+\dot{J}}$, $\mathbf{r}_{p}(f)\in{\mathbb{C}}^{\dot{I}+\dot{J}}$, and $\varepsilon^{\prime}_{p}$ absorbs the incurred modeling error and $\varepsilon_{p}$. Notice that, for simplicity, we use $\dot{I}-1$ past and $\dot{J}$ future taps for both $\overline{\mathbf{S}}_{q}(t,f)$ and $\acute{\mathbf{V}}_{q}(t,f)$, although it might be better to use different taps for different sources. On the other hand, $\mathbf{r}_{p}(f)^{\mathsf{H}}\acute{\mathbf{V}}_{q}(t,f)$ could be a crude approximation of $V_{p}(t,f)$, as there could be multiple directional and diffuse noise sources, rather than a single directional speaker source like in $S_{q}$.
Following (2), the closed-talk mixture is formulated as follows:
$Y_{0}(t,f)$ $=X_{0}(t,f)+V_{0}(t,f)$
$=\mathbf{h}_{0}(f)^{\mathsf{H}}\breve{\mathbf{S}}_{q}(t,f)+V_{0}(t,f)+\varepsilon_{0}(t,f)$
$=\mathbf{h}_{0}(f)^{\mathsf{H}}\breve{\mathbf{S}}_{q}(t,f)+\mathbf{r}_{0}(f)^{\mathsf{H}}\grave{\mathbf{V}}_{q}(t,f)+\varepsilon^{\prime}_{0}(t,f),$ (3)
@@
where we use subscript 00 to denote the close-talk microphone and differentiate it with far-field microphones. In row $2$, $X_{0}(t,f)\approx\mathbf{h}_{0}(f)^{\mathsf{H}}\breve{\mathbf{S}}_{q}(t,f)$ with $\breve{\mathbf{S}}_{q}(t,f)=[S_{q}(t-\ddot{I}+1,f),\dots,S_{q}(t+\ddot{J},f)]\in{\mathbb{C}}^{\ddot{I}+\ddot{J}}$ and $\mathbf{h}_{0}(f)\in{\mathbb{C}}^{\ddot{I}+\ddot{J}}$. In row $3$, $V_{0}(t,f)\approx\mathbf{r}_{0}(f)^{\mathsf{H}}\grave{\mathbf{V}}_{q}(t,f)$, with $\grave{\mathbf{V}}_{0}(t,f)=[V_{q}(t-\ddot{I}+1,f),\dots,V_{q}(t+\ddot{J},f)]\in{\mathbb{C}}^{\ddot{I}+\ddot{J}}$ and $\mathbf{r}_{0}(f)\in{\mathbb{C}}^{\ddot{I}+\ddot{J}}$. In $X_{0}$, the direct-path signal of the target speaker is often much stronger than its reverberation, and hence $X_{0}$ can be largely viewed as the dry source signal. In this case, $\mathbf{h}_{0}(f)$ can be interpreted as a deconvolutional filter that can reverse the time delay and gain decay in the direct-path signal $S_{q}$ to recover the speech source signal.
@@ 3.2 Formulating Speech Enhancement as Blind Deconvolution
As is suggested by UNSSOR [35] and M2M [40], close-talk and far-field mixtures contain weak supervision for speaker separation. One way to exploit the weak supervision for speech enhancement is by solving the following problem, which finds sources, $S_{q}(\cdot,\cdot)$ and $V_{q}(\cdot,\cdot)$, and filters, $\mathbf{g}_{q}(\cdot)$, $\mathbf{h}_{\cdot}(\cdot)$ and $\mathbf{r}_{\cdot}(\cdot)$, that are most consistent with the physical models in (1), (2) and (3):
$\underset{\begin{subarray}{c}S_{q}(\cdot,\cdot),V_{q}(\cdot,\cdot),\\ \mathbf{g}_{q}(\cdot),\mathbf{h}_{\cdot}(\cdot),\mathbf{r}_{\cdot}(\cdot)\end{subarray}}{\text{argmin}}\Big{(}\sum\limits_{t,f}\Big{|}Y_{q}(t,f)-S_{q}(t,f)-\mathbf{g}_{q}(f)^{\mathsf{H}}\widetilde{\mathbf{S}}_{q}(t,f)-V_{q}(t,f)\Big{|}^{2}$
$\quad+\sum\limits_{p=1,p\neq q}^{P}\sum\limits_{t,f}\Big{|}Y_{p}(t,f)-\mathbf{h}_{p}(f)^{\mathsf{H}}\overline{\mathbf{S}}_{q}(t,f)-\mathbf{r}_{p}(f)^{\mathsf{H}}\acute{\mathbf{V}}_{q}(t,f)\Big{|}^{2}$
$\quad+\sum\limits_{t,f}\Big{|}Y_{0}(t,f)-\mathbf{h}_{0}(f)^{\mathsf{H}}\breve{\mathbf{S}}_{q}(t,f)-\mathbf{r}_{0}(f)^{\mathsf{H}}\grave{\mathbf{V}}_{q}(t,f)\Big{|}^{2}\Big{)},$ (4)
where ||$\cdot$|| ⋅ | computes magnitude. This is a blind deconvolution problem [67], which is non-convex and difficult to be solved since the speech source, noise source, and linear filters are all unknown and need to be estimated. It is known not solvable if no prior knowledge is assumed about the sources or filters.
In [35], UNSSOR, which models source priors via unsupervised deep learning, is proposed to tackle this category of blind deconvolution problems. It is shown effective at separating reverberant multi-speaker mixtures to reverberant speaker images in simulated conditions.
3.3 Overview of Adapting M2M for Speech Enhancement
Building upon the preliminary successes of UNSSOR [35] and M2M [40] in speaker separation on simulated data, this paper adapts M2M training for neural speech enhancement and performs training and evaluation on real-recorded data. The high-level idea is to use un- or weakly-supervised deep learning to first estimate the speech and noise sources. With the two sources estimated, filter estimation in (3.2) becomes much simpler linear regression problems, where closed-form solutions exist and can be readily computed. With the sources and filters estimated, we can then compute a loss defined similarly to the objective in (3.2) to regularize the two source estimates to have them respectively approximate the speech and noise sources.
@@ 3.3 Overview of Adapting M2M for Speech Enhancement
Figure 2: Illustration of SuperM2M, which consists of (a) M2M training on real close-talk and far-field mixture pairs (described in first paragraph of Section 4); and (b) supervised training on simulated far-field mixtures (described in Section 5). M2M trains the DNN to separate far-field mixtures to two sources that can be linearly filtered to reconstruct far-field and close-talk mixtures. In SuperM2M, we alternately feed in mini-batches of real mixtures and mini-batches of simulated mixtures, and train the same DNN by alternating between M2M training on real data and supervised learning on simulated data.
4 M2M for Speech Enhancement

### v3

# SuperM2M: Supervised and Mixture-to-Mixture Co-Learning for Speech Enhancement and Noise-Robust ASR

@@ Abstract
Neural speech enhancement , noise-robust ASR
journal: Neural Networks
@@ 1 Introduction
On the other hand, during data collection, besides using far-field microphones to record target speech, a close-talk microphone is often placed near the target speaker to collect its close-talk speech (e.g., in the CHiME [36], AMI [37], AliMeeting [38], and MISP [39] setup). [^1] Close-talk speech is almost always recorded together with far-field speech in speech separation and recognition datasets, as it is much easier for humans to annotate word transcriptions and speaker activities based on close-talk recordings (where the target speech is very strong) than far-field recordings. Although the close-talk microphone can also pick up non-target signals, the recorded close-talk mixture typically has a much higher signal-to-noise ratio (SNR) of the target speaker than any far-field mixtures. In our recent mixture to mixture (M2M) algorithm [40], which builds upon UNSSOR, we find that, besides far-field mixtures, the close-talk mixture can also be leveraged as a weak supervision for training DNNs to separate mixed speakers.
@@
In this context, we propose to extend UNSSOR [35] and M2M [40] for speech enhancement and robust ASR, where we train and evaluate the proposed algorithms not only on simulated data but also on real data. [^2] This long-form paper is a journal expansion of our preliminary work in UNSSOR [35] and M2M [40]. Our major goal is to show whether the resulting algorithms can yield better generalization to real data than purely-supervised models trained on simulated data, a demonstration that is missing in UNSSOR [35] and M2M [40]. Without this demonstration, the evidence supporting whether this un- and weakly-supervised line of research is worth investigating would be lacking, especially considering the dominance and simplicity of purely-supervised approaches based on simulated training data. We summarize the key contributions of this paper as follows:
@@
  • Considering noise sources as a single, combined source, we formulate the training of speech enhancement models on real data as solving a blind deconvolution problem, following the formulations in UNSSOR [35] and M2M [40] designed for speaker separation.
  • To account for the case when close-talk mixture is not time-synchronized with its paired far-field mixtures, we propose an on-the-fly filter tap estimation algorithm that can deal with the time-synchronization issue.
  • We propose SuperM2M, a co-learning strategy which trains the same DNN model by alternating between M2M training on real data and supervised learning on simulated data. This way, M2M can benefit from massive simulated training data, especially when the real training data is scarce. In addition, we find that this strategy can help mitigate the weaknesses of UNSSOR and M2M on source permutation, frequency permutation, and source ambiguity (all of which will be detailed in Section 6.1).
@@
The evaluation results provide an evidence supporting the strong potential of our un- and weakly-supervised line of research for speech enhancement. A sound demo is provided in the link below. [^3] https://zqwang7.github.io/demos/SuperM2M_demo/index.html
@@ 2.1 Frontend Enhancement and Robust ASR
Differently, in this paper we aim at building neural speech enhancement models whose immediate estimate itself can have low distortion to target speech and high reduction to non-target signals, especially on real test data. We find that, on the challenging real test data of CHiME-4 [42], the immediate output of SuperM2M bears low distortion to target speech and high reduction to non-target signals, and feeding it directly to strong ASR models for recognition yields strong performance.
@@ 3 Problem Formulation
We start with describing the hypothesized physical models, then propose to formulate speech enhancement as a blind deconvolution problem, and at last overview the proposed algorithm for speech enhancement.
@@ 3.1 Physical Model
In noisy-reverberant conditions with a compact far-field $P$-microphone array and a single target speaker wearing a close-talk microphone, the physical model for each real-recorded mixture can be formulated, in the short-time Fourier transform (STFT) domain, as follows.
At a designated reference far-field microphone $q\in\{1,\dots,P\}$, the real-recorded mixture is formulated as
$Y_{q}(t,f)$ $=X_{q}(t,f)+V_{q}(t,f),$ (1)
where $t$ indexes $T$ frames, $f$ indexes $F$ frequency bins, and $Y_{q}(t,f)$, $X_{q}(t,f)$, and $V_{q}(t,f)$ respectively denote the STFT coefficients of the mixture, speaker image of the target speaker, and non-speech signals at time $t$, frequency $f$ and microphone $q$. In the rest of this paper, when dropping indices $t$ and $f$, we refer to the corresponding spectrograms. We emphasize that $V_{q}$ could contain multiple strong, non-stationary directional as well as diffuse noises. In this paper, we model them using a single, combined source.
Let us denote the index of the close-talk microphone as 00, which is different from the set of indices $\{1,\dots,P\}$ used for far-field microphones. This way, we can index any of the non-reference microphones [^4] By non-reference microphones, we mean the close-talk microphone and $P-1$ non-reference far-field microphones. using $p\in\{0,1,\dots,P\}$, with $p\neq q$. We then formulate the mixture captured at microphone $p$ as
@@
$=\mathbf{h}_{p}(f)^{\mathsf{H}}\overline{\mathbf{X}}_{q}(t,f;p)+V_{p}(t,f)+\varepsilon_{p}(t,f)$
$=\mathbf{h}_{p}(f)^{\mathsf{H}}\overline{\mathbf{X}}_{q}(t,f;p)+\mathbf{r}_{p}(f)^{\mathsf{H}}\overline{\mathbf{V}}_{q}(t,f;p)+\varepsilon^{\prime}_{p}(t,f).$ (2)
@@
In row $2$, following narrowband approximation [66, 51], we approximate the speaker image captured at microphone $p$ (i.e., $X_{p}(\cdot,f)$) as a linear convolution between the speaker image captured at microphone $q$ (i.e., $X_{q}(\cdot,f)$) and a linear filter $\mathbf{h}_{p}(f)$. That is, $X_{p}(t,f)\approx\mathbf{h}_{p}(f)^{\mathsf{H}}\overline{\mathbf{X}}_{q}(t,f;p)$, where $\overline{\mathbf{X}}_{q}(t,f;p)=[X_{q}(t-I_{p}+1,f),\dots,X_{q}(t+J_{p},f)]\in{\mathbb{C}}^{I_{p}+J_{p}}$ and $\mathbf{h}_{p}(f)\in{\mathbb{C}}^{I_{p}+J_{p}}$, and we use $\varepsilon_{p}$ to denote the modeling error. Notice that the past and future filter taps, $I_{p}$ and $J_{p}$, are dependent on the microphone index (i.e., $p$) of the speaker image $X_{p}$ we want to approximate. $\mathbf{h}_{p}(f)$ can be interpreted as the relative transfer function (RTF) relating the speaker image $X_{q}$ (captured by the reference far-field microphone $q$) to the speaker image captured at another microphone $p$ (i.e., $X_{p}$). In row $3$, we use the same trick to approximate non-speech signals $V_{p}(t,f)\approx\mathbf{r}_{p}(f)^{\mathsf{H}}\overline{\mathbf{V}}_{q}(t,f;p)$, where $\overline{\mathbf{V}}_{q}(t,f;p)=[V_{q}(t-I_{p}+1,f),\dots,V_{q}(t+J_{p},f)]\in{\mathbb{C}}^{I_{p}+J_{p}}$ and $\mathbf{r}_{p}(f)\in{\mathbb{C}}^{I_{p}+J_{p}}$, and $\varepsilon^{\prime}_{p}$ absorbs the incurred modeling error and $\varepsilon_{p}$.
To facilitate understanding, we provide several comments:
  • For simplicity, we use $I_{p}-1$ past and $J_{p}$ future taps for both $\overline{\mathbf{X}}_{q}(t,f;p)$ and $\overline{\mathbf{V}}_{q}(t,f;p)$, although it might be better to use different taps for different sources.
  • For the filter taps $I_{p}$ and $J_{p}$, we have a subscript $p$ to indicate that we can use different number of filter taps for each non-reference microphone $p$. It may be a good idea to use different filter taps for the non-reference far-field microphones and the close-talk microphone, especially if they are placed on different devices or not synchronized with each other.
  • $\mathbf{r}_{p}(f)^{\mathsf{H}}\overline{\mathbf{V}}_{q}(t,f;p)$ could be a crude approximation of $V_{p}(t,f)$, as there could be multiple directional and diffuse noise sources in $V$, rather than a single directional source like in $X_{q}$.
@@
  • In the speaker image captured by the close-talk microphone (i.e., $X_{0}$), the direct-path signal of the target speaker is often much stronger than its reverberation. Therefore, $X_{0}$ can be largely viewed as the dry source signal. In this case, $\mathbf{h}_{0}(f)$ can be interpreted as a deconvolutional filter that can reverse the speaker image $X_{q}$ back to the speech source signal.
@@ 3.2 Formulating Speech Enhancement as Blind Deconvolution
As is suggested by our preliminary studies, UNSSOR [35] and M2M [40], both close-talk and far-field mixtures contain weak supervision for speaker separation. They [35, 40] exploit the weak supervision by formulating speaker separation as solving a blind deconvolution problem.
@@
Figure 1: Illustration of SuperM2M, which consists of (a) M2M training on real close-talk and far-field mixture pairs (described in Section 4); and (b) supervised training on simulated far-field mixtures (described in Section 6). M2M trains the DNN to separate far-field mixtures to two sources that can be linearly filtered to reconstruct far-field and close-talk mixtures. In SuperM2M, we alternately feed in mini-batches of real mixtures and mini-batches of simulated mixtures, and train the same DNN by alternating between M2M training (if the input mixtures are real) and supervised learning (if the input mixtures are simulated).
Similar to speaker separation, speech enhancement can be viewed as two-source separation, with one source being speech and the other being all non-speech sources combined. With this understanding, following UNSSOR [35] and M2M [40], we can also formulate speech enhancement as solving a blind deconvolution problem, and adapt UNSSOR and M2M for speech enhancement. In detail, the problem, formulated below, finds sources, $X_{q}(\cdot,\cdot)$ and $V_{q}(\cdot,\cdot)$, and filters, $\mathbf{h}_{\cdot}(\cdot)$ and $\mathbf{r}_{\cdot}(\cdot)$, that are most consistent with the physical models in (1) and (2):
$\underset{\begin{subarray}{c}X_{q}(\cdot,\cdot),V_{q}(\cdot,\cdot),\\ \mathbf{h}_{\cdot}(\cdot),\mathbf{r}_{\cdot}(\cdot)\end{subarray}}{\text{argmin}}\Big{(}\sum\limits_{t,f}\Big{|}Y_{q}(t,f)-X_{q}(t,f)-V_{q}(t,f)\Big{|}^{2}+$
$\sum\limits_{p=0,p\neq q}^{P}\sum\limits_{t,f}\Big{|}Y_{p}(t,f)-\mathbf{h}_{p}(f)^{\mathsf{H}}\overline{\mathbf{X}}_{q}(t,f;p)-\mathbf{r}_{p}(f)^{\mathsf{H}}\overline{\mathbf{V}}_{q}(t,f;p)\Big{|}^{2}\Big{)},$ (3)
where ||$\cdot$|| ⋅ | computes magnitude. As is suggested in UNSSOR [35], this problem is non-convex and difficult to be solved, since the speech source, noise source, and linear filters are all unknown and need to be estimated. It is known not solvable if no prior knowledge is assumed about the sources or filters [67]. In our preliminary work, UNSSOR [35], which models source priors via unsupervised deep learning, is proposed to tackle this category of blind deconvolution problems. It is shown effective at separating reverberant multi-speaker mixtures to reverberant speaker images in simulated conditions.
3.3 Overview of Proposed Algorithms
Building upon the preliminary successes of UNSSOR [35] and M2M [40] in speaker separation on simulated data, this paper adapts UNSSOR and M2M training for neural speech enhancement and performs training and evaluation on real-recorded data.
We first give UNSSOR and M2M a review in Section 4. We then propose to improve them by addressing potential synchronization issues between close-talk and far-field microphones in Section 5. Next, in Section 6 we point out several weaknesses of UNSSOR and M2M, and propose to address them by combining UNSSOR and M2M training on real data with supervised learning on simulated data.
4 M2M Review, and Adapting It for Speech Enhancement
Since we formulate speech enhancement as two-source separation, we can readily adapt UNSSOR and M2M (originally designed for speaker separation) for speech enhancement. This section reviews UNSSOR and M2M training, but in the context of adapting them for speech enhancement.